OpenStudy (anonymous):

Because not all airline passengers show up for their reserved seat, an airline sells 75 tickets for a flight that holds only 72 passengers. The probability that a passenger does not show up is 0.10, and the passengers behave independently. What is the probability that more than 72 passengers will show up? What are the mean and standard deviation of the number of passengers who show up?

OpenStudy (anonymous):

Anyone know how do to do this?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

your problem follows a binomial distribution.

OpenStudy (anonymous):

could you possibly do that out? i got an answer of .9582 for p and a mean of 67.5 and a std of 2.598 i just want to verify those answers

OpenStudy (zarkon):

I get the same mean as you, but that is it.

OpenStudy (anonymous):

okay, how did u do the rest? and what did u get?

OpenStudy (zarkon):

P(more than 72 passengers will show up) =P(73 or 74 or 75 passengers show up) =P(73 passengers show up)+P(74 passengers show up)+P(75 passengers show up)

OpenStudy (zarkon):

\[=\sum_{k=73}^{75}{75\choose k}(.9)^k(.1)^{75-k}\]

OpenStudy (zarkon):

\[\approx 0.016128757292\]

OpenStudy (anonymous):

would the method you used, be used in an entry level stats class?

OpenStudy (zarkon):

I don't see why not.

OpenStudy (anonymous):

The method i used was p(z<= (x-u)/(up))

OpenStudy (zarkon):

most intro stat classes go over the binomial distribution (including the density function)

OpenStudy (zarkon):

you are doing a normal approximation?

OpenStudy (anonymous):

i guess, i dont really know

OpenStudy (zarkon):

for this problem we have n=75 and p=.9 so our mean is 75*.9=67.5 and the sd is \[\sigma=\sqrt{np(1-p)}=\sqrt{75\cdot.9\cdot.1}=2.598\] so i will agree with you there

OpenStudy (zarkon):

have you done any continuity corrections in class?

OpenStudy (zarkon):

or do you not know what I am talking about

OpenStudy (anonymous):

we have not

OpenStudy (zarkon):

if you don't use a continuity correction you should get .041632219204 (approx 1 minus the answer you gave above)

OpenStudy (zarkon):

did you compute P(Z>1.732)?

OpenStudy (anonymous):

Thats where i went wrong. i used the wrong symbol

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (zarkon):

np

OpenStudy (anonymous):

do you have time for another problem?

OpenStudy (zarkon):

maybe

OpenStudy (anonymous):

Supposed the contamination particle size (in micrometers) can be modeled as f(x)=3x^-4 for x>1 and f(x)=0 and for x\[\le\]1

OpenStudy (anonymous):

a) confirm that f(x) is a probability density function b) give the cumulative distribution function c)determine the mean d) what is the probability that the size of a random particle will be less than 5 micrometers?

OpenStudy (zarkon):

I assume you know some calculus.

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

you need to show that \[\int\limits_{1}^{\infty}3x^{-4}dx=1\]

OpenStudy (anonymous):

i have some answers, but i have to sign off. if you could answer this i would greatly appreciate it!

OpenStudy (zarkon):

CDF \[F(x)=\left\{\begin{matrix}0 & \text{if }x<1 \\ \frac{x^3-1}{x^3} & x\ge1\end{matrix}\right.\]

OpenStudy (zarkon):

mean is 3/2

OpenStudy (zarkon):

prob less than 5 is \[\frac{124}{125}\]